### v3

# references.bib

Abstract
The Backpropagation algorithm has often been criticised for its lack of biological realism. In an attempt to find a more biologically plausible alternative, the recently introduced Forward-Forward algorithm replaces the forward and backward passes of Backpropagation with two forward passes. In this work, we show that the internal representations obtained by the Forward-Forward algorithm can organise into category-specific ensembles exhibiting high sparsity – composed of a low number of active units. This situation is reminiscent of what has been observed in cortical sensory areas, where neuronal ensembles are suggested to serve as the functional building blocks for perception and action. Interestingly, while this sparse pattern does not typically arise in models trained with standard Backpropagation, it can emerge in networks trained with Backpropagation on the same objective proposed for the Forward-Forward algorithm. These results suggest that the learning procedure proposed by Forward-Forward may be superior to Backpropagation in modelling learning in the cortex, even when a backward pass is used.
1 Introduction
Deep Learning is a highly effective approach to artificial intelligence, with tremendous implications for science, technology, culture, and society. At its core, there is the Backpropagation (Backprop) algorithm [rumelhart1986learning], which efficiently computes the gradients necessary to optimise the learnable parameters of an artificial neural network. Backprop, however, lacks biological plausibility [stork1989backpropagation] – leading to many attempts to address the issue. The most recent of such, the Forward-Forward algorithm [hinton2022forward], eliminates the need to store neural activities and propagate error derivatives along the network.
In a standard classification context, the application of Forward-Forward requires the designation of positive and negative data. For example, to classify images, one could assign positive (or negative) data to those images having their correct (or incorrect, respectively) class embedded via one-hot encoding at the border (as shown in Figure 1, Panel A). The Forward-Forward algorithm then learns to discriminate between positive and negative data by optimising a goodness function (e.g., the $\ell_{2}$ norm of the activations), akin to contrastive learning [chen2020simple]. Satisfactory results have been observed [hinton2022forward] for classification tasks on Mnist [lecun2010mnist], a well-known benchmark dataset.
This work goes beyond mere performance, investigating the presence of similarities between biological and artificial neuronal ensembles. Our experiments demonstrate sparsity in representations learned using Forward-Forward, and similarities between these and cortical ensembles found in early stages of sensory processing [miller2014visual]. Neurons that form ensembles are highly specialised, and are found to activate selectively when presented with positive data, while only exhibiting minimal activation to negative data or unstructured stimuli. Our results also indicate that in such representations there may exist collective suppression mechanisms similar to those of biological inhibitory neurons [yuste2015neuron].
Finally, though optimising the cross-entropy loss for the same classification task does not appear to produce the sparse ensembles we observe, the phenomenon may not solely be due to the use of the Forward-Forward algorithm. In fact, similar results are obtained by optimising the same goodness function of Forward-Forward, with Backprop instead. This suggests that more focus should be put on the purpose and biological meaning of the loss function rather than the training algorithm [richards2019deep].
Figure 1: Activation patterns in a Multi-Layer Perceptron trained with the Forward-Forward algorithm, on the Mnist dataset.
Panel A Examples of activation patterns in response to a typical input. Images show the activation value for network units, arranged as a matrix only for the sake of clarity; darker squares represent more active neurons.
Panel B Activation value of each neuron in the first hidden layer (Layer 1), averaged on all images of a given class. Neuron index on the $x$ axis; average activation on the $y$ axis. Blue dots indicate units that are considered active according to the leave-one-out (LOO) method described in subsection 3.5.
Panel C Activation map for neurons in Layer 1 for all images, grouped by class. A blue dot in position $(x,y)$ indicates that neuron $x$ is activated by input $y$; colorscale represents the intensity of such activation. Horizontal bands mark different categories; blue vertical stripes mark active, category-specific neurons. Each input category activates consistently a specific sets of neurons (ensemble).
2 Related Work
In the section that follows, we summarise key aspects of the Forward-Forward algorithm and the main findings pertaining identification and characterisation of biological neuronal ensembles in the cortex.
2.1 Forward-Forward
The Forward-Forward algorithm [hinton2022forward] is a recently proposed learning algorithm for artificial neural networks, whose main premise is the ability to overcome the notorious biological implausibility of Backprop [rumelhart1986learning]. In fact, while the effectiveness of Backprop makes it the standard algorithm for training neural networks, it is based on biologically unrealistic assumptions, such as the need to propagate information forwards and backwards through the network [richards2019deep].
Forward-Forward owes its name to the fact that it replaces the backward pass with an additional forward pass. The two forward passes are executed on different data, named positive and negative data. During training, the objective of Forward-Forward is to maximise a so-called goodness function of the neural activations (e.g. the $\ell_{p}$ norm) on positive data and minimise it on negative data. In a simple image classification setting, such as the one we adopt in this paper, one could encode a class label at the border of images, by one-hot encoding it with a white pixel (as shown in Figure 1, Panel A). Then, following the definition from [hinton2022forward], positive data are those for which the encoded label matches the ground truth label, while the opposite holds for negative data.
Layers are trained separately and sequentially, and learn to discriminate between positive and negative data by maximising and minimising their goodness, according to the data presented. Crucially, activations are normalised before being passed to the subsequent layer, to prevent layers from relying on the goodness computed by their predecessors. From the biological point of view, normalisation is known to be a “canonical neural computation” [carandini2012]. At test time, when a new unlabelled sample has to be categorised, many copies of the image are created, each with a different one-hot encoded label. These are then fed into the neural network to obtain a goodness score. Finally, the image gets classified in the category that produced the maximum goodness value.
In the seminal Forward-Forward paper [hinton2022forward], satisfactory classification results are reported on the standard handwritten digit recognition dataset Mnist, with the definition of positive and negative data described above, and using the $\ell_{2}$ or $\ell_{1}$ norm of activations as goodness function.
In a recent theoretical work, it has been analytically shown that under somewhat mild assumptions sparsity emerges in Forward-Forward layers [yang2023theory] as a consequence of optimizing the Forward-Forward loss. While these theoretical results are derived for a single layer, they bolster the reliability of our experimental findings.
2.2 Neuronal ensembles
In Neuroscience, ensembles are defined as sparse groups of neurons that co-activate during spontaneous activity or in response to sensory stimuli. Remarkably, ensembles – rather than single neurons – have long been suggested to be emergent functional blocks of cortical activity [yuste2015neuron, hebb2005organization, harris2005neural, buzsaki2010neural] and have a prominent role in sensory processing, memory [hopfield1982neural] and behaviour [carrillo2019controlling]. As shown by [miller2014visual], during visual processing the spiking activity in the cortex is dominated by ensembles, whose properties cannot be accounted for by the independent firing activity of neurons in isolation. Additionally, ensembles can be both generated by sensory stimuli (visual, in this case) or by the spontaneous activity of the network. It is also shown that single neurons can participate in more than one ensemble, thus maximising the encoding potential of the network. Interestingly, the same ensembles are evoked both when reacting to stimuli and during spontaneous activity, indicating that they may serve as inherent fundamental components giving rise to visual responses. Findings from [yoshida2020natural] confirm that visual inputs activate sparse ensembles in the primary visual cortex (V1). Moreover, images can be decoded starting from a low number of highly responsive neurons (i.e. a highly sparse ensemble), and adding more does not increase (or even harms) decoding performance. These neurons consistently and reliably encode the same visual inputs across multiple trials. For that reason, partially overlapping receptive fields may play a role in achieving sparse and robust representation of information. Consequently, collecting the activity of highly responsive neurons emerges as an optimal decoding strategy for downstream neurons. The perdurance of ensembles – they have the capacity to endure for weeks [perez2021long] – suggests that they could also potentially serve as a foundation for the long-term representation of perceptual states or memories.
Finally, the idea of neuronal ensembles has inspired the development of computational models [papadimitriou2020brain]. E.g., [lewicki2004sparse] shows that sparse and redundant (hence more robust) representations are optimal for encoding natural images with noisy or unreliable neurons, a finding as also confirmed by [field1994goal] and [olshausen2004sparse].
3 Methods
In this work, we investigate and compare the representations produced by three models:
  • A classifier in the style of that used by [hinton2022forward], trained with Forward-Forward (FF);
  • A classifier identical to the above, but trained end-to-end with Backprop to optimise the same goodness function (BP/FF);
  • A classifier trained with Backprop on the categorical cross-entropy loss, as customary (BP).
Such different scenarios are further analysed individually in subsection 3.2, subsection 3.3 and subsection 3.4, respectively.
3.1 Data
The datasets we use to train and test the models described so far are Mnist [lecun2010mnist], FashionMnist [xiao2017fashion], Svhn [netzer2011reading] and Cifar10 [alex2009learning]. Details on these datasets are provided in the Appendix B.
3.2 Model trained with Forward-Forward (FF)
Our FF model is inspired by the architecture proposed by [hinton2022forward] – and likewise trained according to the Forward-Forward algorithm. It consists of three fully-connected layers, each composed by $1000$ units in the case of Mnist and FashionMnist, and $3072$ units in the case of Svhn and Cifar10. Each linear layer is followed by elementwise ReLU non-linearities. Both during training and inference, the layer-wise $\ell_{2}$ norm is used as the goodness function of choice; correspondingly, $\ell_{2}$ normalisation is performed between subsequent layers.
To define positive and negative data, a one-hot-encoded class vector is embedded at the top-left corner of images. Prior to such embedding, these pixels are set to black colour. Then, in the case of positive data, the pixel corresponding to the true class is switched to the maximum value elsewhere observed in the image, while in the case of negative examples such value is randomly assigned to one of the other pixels of the embedding vector.
During training, the weights are optimised by minimising the loss function $L=\log(1+e^{G_{neg}-G_{pos}})$, where $G_{neg}$ and $G_{pos}$ are, respectively, the goodness value for negative and positive data. At inference time, for each layer, the goodness values corresponding to every possible label are converted into a probability using softmax. By performing this step for each layer, they can contribute equally to the prediction.
3.3 Model trained with Backpropagation on the goodness objective (BP/FF)
The architecture of the FF model, while designed to be optimised using the Forward-Forward algorithm, can be trained seamlessly with Backpropagation on the same goodness maximisation/minimisation objective. Indeed, keeping the definition of positive and negative data introduced for FF, one could simply use Backprop to optimise the goodness-based loss from the Forward-Forward algorithm.
In detail, positive and negative data are fed to the network during the forward step, and the overall goodness of the internal representation is evaluated. The backward pass is then executed, and parameters are optimised to achieve the same goal as the FF model. It is worth pointing out that, in this case, the goodness is maximised globally instead of layer-by-layer (i.e., locally).
3.4 Model trained with Backpropagation on the cross-entropy loss (BP)
The FF and BP/FF models are also compared to a standard neural classifier, serving as a baseline. For such purpose, a multi-layer perceptron is employed. The model shares the same number of layers, layerwise neuron count, and non-linear activation function choice with FF and BP/FF. The only architectural difference between the BP model and the other two is the addition of a final softmax layer, to suitably shape and scale the output for the classification task. The model is trained end-to-end with Backprop on the categorical cross-entropy loss.
3.5 Analysis of representations
For each model described, we analyse the internal representation emerging at each layer. We limit our analysis to data belonging to the test set (i.e. not seen during training) and correctly classified by the respective model. Concretely, the representation of a single image is a $n$-dimensional vector composed by the activations (after the ReLU non-linearity) of all the units in the layer. For each layer, we extract a representation matrix $X$ of size $(M,n)$, where $M$ is the total number of test images (correctly classified) and $n$ is the number of neurons in the layer considered.
Sparsity
For each representation vector $x$ we assign a sparsity measure following the notion of sparsity introduced in [hoyer2004non]:
$S(x)=\frac{\sqrt{n}-\frac{\lVert x\rVert_{1}}{\lVert x\rVert_{2}}}{\sqrt{n}-1}$
With this definition, when $S(x)=1$ the vector $x$ contains only one non-zero component representing the case of an extreme sparsity. The other limiting case is the one in which all the components of $x$ are equal in magnitude, in this case $S(x)=0$. The sparsity function $S$ interpolates smoothly between these two extremes. The sparsity of a layer representation is obtained by averaging the sparsity of its component vectors $S=\frac{1}{M}\sum_{i=1}^{M}S(x_{i})$.
Ensembles
To detect the emergence of category-specific ensembles, within each model and dataset combination, we adopt the following method. The idea is that a neuron should be considered active and part of an ensemble if it activates consistently and selectively when the network receives input data that belongs to that category.
We start by defining a category-specific representation matrix $X_{c}$, of shape $(M_{c},n)$, where $M_{c}$ is the number of correctly classified test images of the given category. Then, we compute the average activation of each hidden unit across all samples: $\overline{x_{j}}=\frac{1}{M_{c}}\sum_{i=1}^{M_{c}}(X_{c})_{ij}$; and the leave-one-out average of the averages $\text{LOO}_{j}=\frac{1}{n-1}\sum_{i\neq j}\overline{x_{j}}$. We then classify a neuron $i$ as active (i.e. part of an ensemble) if $\overline{x_{i}}>\text{LOO}_{i}$. Examples of average activation profiles and of ensembles are reported in Figure 1.
The output of the ensemble computation is a set of active units for each category: $\mathcal{E}^{c}=\left\{e^{c}_{1},e^{c}_{2},\dots e^{c}_{n_{c}}\right\},\forall c\in\{1,2,\dots,C\}$, where $n_{c}$ is the number of active units for category $c$. Once the ensembles are defined, it is possible to look at units that are shared across categories $c$ and $c^{\prime}$ by considering $\mathcal{E}^{c}\cap\mathcal{E}^{c^{\prime}}$. The size of the shared units is naturally measured by $\mid\mathcal{E}^{c}\cap\mathcal{E}^{c^{\prime}}\mid$.
We can also measure the similarity between two ensembles using the Jaccard similarity index (intersection over union): $J(\mathcal{E}^{c},\mathcal{E}^{c^{\prime}})=\frac{\mid\mathcal{E}^{c}\cap\mathcal{E}^{c^{\prime}}\mid}{\mid\mathcal{E}^{c}\cup\mathcal{E}^{c^{\prime}}\mid}$. Examples of shared units are reported in Figure 3.
When the sparsity $S$ of a representation is low, ensembles are typically ill-defined as too many neurons are significantly active simultaneously and the notion of active unit tend to blur. To set a threshold, we will consider values of $S$ below $0.5$ as not-sparse, and in these cases, we do not define ensembles out of the representation.
4 Results
We describe our findings for the three models introduced, on the Mnist, FashionMnist, Svhn and Cifar10 datasets. In particular, we focus on the unique properties of representations obtained within the FF model, i.e. a model trained with Forward-Forward on its natural goodness objective. Such properties – e.g. the emergence of category-specific ensembles and the presence of shared units across them – establish a link between neural networks trained with the Forward-Forward algorithm and biological cortical networks described in subsection 2.2.
4.1 Classification accuracy
Before delving into the main results of this work, we evaluate the performances of our models on the classification tasks at hand. Table 1 contains results in terms of test set classification accuracy for all models we employed – FF, BP/FF and BP – on Mnist, FashionMnist, Svhn and Cifar10. While some of these accuracy values are far from the state-of-the-art (i.e., respectively, $0.997$ [ciresan2010bigdeep], $0.931$ [xiao2017fashion], $0.860$ [pitsios2017] and approximately $0.7$ [lin2015far], for fully-connected networks), they are a solid ground on which to build our subsequent investigations. Training details and hyperparameters for all models are reported in Appendix C.
Table 1: Test-set classification accuracy for the models considered in our investigation. Results expressed as mean $\pm$ std. dev. over 10 runs with independent randomised weight initialisation.
Dataset FF BP/FF BP
Mnist $0.94\pm 0.008$ $0.969\pm 0.001$ $0.982\pm 0.001$
FashionMnist $0.849\pm 0.002$ $0.877\pm 0.002$ $0.892\pm 0.004$
Svhn $0.716\pm 0.002$ $0.799\pm 0.004$ $0.793\pm 0.145$
Cifar10 $0.484\pm 0.004$ $0.521\pm 0.006$ $0.564\pm 0.004$
4.2 Forward-Forward elicits sparse neuronal ensembles
Figure 2: Sparsity of category-specific representations. We report the sparsity of representations - computed as described in subsection 3.5 - for the three models FF, BP/FF and BP on the Mnist dataset (same models as in the first row of Table 1). Sparsity values are the average on 10 runs.
The FF and BP/FF models – based on the original Forward-Forward network architecture, and trained according to the goodness objective (subsection 3.2 and subsection 3.3) – exhibit typically high sparsity levels in their representations, in clear contrast with BP (see Figure 2 and Table 2).
Table 2: Average sparsity for all combinations of model, dataset and layer, according to the definition given in subsection 3.5. Results are expressed as mean $\pm$ std. dev. computed over 10 runs with independent random weights initialization.
Model Layer Mnist FashionMnist Svhn CIFAR10
1 $0.922\pm 0.001$ $0.85\pm 0.002$ $0.83\pm 0.001$ $0.77\pm 0.001$
FF 2 $0.813\pm 0.019$ $0.605\pm 0.015$ $0.706\pm 0.001$ $0.728\pm 0.002$
3 $0.618\pm 0.074$ $0.628\pm 0.013$ $0.489\pm 0.004$ $0.566\pm 0.002$
1 $0.895\pm 0.005$ $0.81\pm 0.007$ $0.783\pm 0.003$ $0.753\pm 0.004$
BP/FF 2 $0.747\pm 0.013$ $0.851\pm 0.007$ $0.95\pm 0.003$ $0.932\pm 0.003$
3 $0.131\pm 0.011$ $0.065\pm 0.009$ $0.133\pm 0.011$ $0.135\pm 0.009$
1 $0.315\pm 0.003$ $0.352\pm 0.003$ $0.47\pm 0.02$ $0.478\pm 0.016$
BP 2 $0.193\pm 0.004$ $0.241\pm 0.005$ $0.524\pm 0.212$ $0.3\pm 0.18$
3 $0.225\pm 0.006$ $0.248\pm 0.006$ $0.232\pm 0.106$ $0.164\pm 0.006$
When the sparsity level is sufficiently high (S > 0.5) we are able to identify small sets of neurons (ensembles) that consistently co-activate across all the samples of the same class, similarly to what has been observed in cortical representations [yuste2015neuron, miller2014visual, harris2005neural].
Figure 1 (Panels B, C) shows an example of average neuron activations for each class in Layer 1 of the FF model trained on Mnist, and showcases the emergence of sparse, category-specific, ensembles (see the Appendix D for a similar visualisation for Layers 2 and 3 of the same model and Appendix E for Layer 1 in all the models).
These representations often elicit a remarkably low fraction of active units: ensembles composed by a few percent of the neurons in a layer are typically observed both for the simplest datasets (Mnist, FashionMnist) and for more complex datasets (Svhn, Cifar10), with a slight tendency of the FF model to create larger ensembles in the latter case (Table 3).
[…]
Model Layer Mnist FashionMnist Svhn CIFAR10
1 $3.69\pm 0.09$ $5.02\pm 0.14$ $10.3\pm 0.15$ $16.08\pm 0.09$
FF 2 $5.31\pm 0.35$ $18.46\pm 0.66$ $21.28\pm 0.23$ $21.2\pm 0.3$
3 $1.36\pm 0.36$ $20.59\pm 0.63$ $4.48\pm 0.52$ $4.86\pm 0.51$
1 $8.58\pm 0.23$ $13.24\pm 0.31$ $15.07\pm 0.16$ $13.3\pm 0.13$
BP/FF 2 $13.18\pm 0.67$ $8.45\pm 0.47$ $5.08\pm 0.19$ $5.55\pm 0.28$
3 - - - -
4.3 Semantically similar classes elicit ensembles with shared neurons
Drawing a parallel with a phenomenon observed in Neuroscience [yoshida2020natural], related categories can be expected to share units of their ensembles. This is indeed what we observe, as shown in Figure 3. Results are reported for FashionMnist, where different classes of clothes or shoes may contain a common share of visual features. In this regard, we observe a clear tendency to share units between similar classes – e.g. across representations of pullover, coat and shirt.
In the following section, we provide evidence that a unit can be shared across two ensembles even if one of these refers to an unseen category (i.e., excluded from the training set but whose representation, extracted at test time, generates a valid ensemble), as we show in Figure 4 (Panel C).
Figure 3: Semantically similar classes in FashionMnist can elicit ensembles with shared neurons.
Panel A The ensembles elicited in the first hidden layer of FF by two example inputs. Red circles indicate the active units which are shared between the two categories.
Panel B Element $i,j$ of the matrix indicates how many units are shared between the ensembles of category $i$ and category $j$ (normalized by the ensemble sizes), by using the Jaccard similarity index: $J(\mathcal{E}^{i},\mathcal{E}^{j})=\frac{\mid\mathcal{E}^{i}\cap\mathcal{E}^{j}\mid}{\mid\mathcal{E}^{i}\cup\mathcal{E}^{j}\mid}$. The results are referred to a single training run.
4.4 Representations of unseen categories can elicit well-defined ensembles
We investigate the ability of a trained FF model to respond to unseen categories with a coherent activation pattern which is typical of the ensembles we found on the categories seen at training time. To this end, we repeatedly train FF on FashionMnist, removing one category at a time. Then, we extract the representation of the missing category, and verify if a valid ensemble is formed. We find that in all the ten cases, a new ensemble is formed, sharing the same characteristics of the ones emerging for seen categories, apparently with the only exception of a lower average activation of their constituent units (see Figure 4 for one example, and the Appendix F for a more detailed account).
In several cases, we also found that the ensembles of unseen categories share units with the ensembles of seen categories, when endowed with similar visual features or semantics (Figure 4, Panel C). A more extensive exploration of these cases is also reported in the Appendix F.
Figure 4: The representation of unseen categories forms a valid ensemble in FF trained on FashionMnist.
Panel A Activation patterns in response to the different categories in the first hidden layer. The unseen category (Sandal), surrounded by red lines, produces a relatively weaker but well-defined ensemble-like activation pattern.
Panel B Activation value of each neuron, averaged on all images of the unseen category. Neuron index on the $x$ axis; average activation on the $y$ axis. Blue dots indicate units that are considered active according to the method described in subsection 3.5.
Panel C Ensembles of unseen categories can share units with the ensembles of the other categories. Element $i,j$ of the matrix indicates how many units are shared between the ensembles of category $i$ and category $j$: $\mid\mathcal{E}^{i}\cap\mathcal{E}^{j}\mid$. The results are referred to a single training run.
5 Discussion and conclusions
The main finding of our work is that artificial neural networks – trained with the Forward-Forward algorithm – can elicit biologically plausible representations in the form of sparse neuronal ensembles [yuste2015neuron, miller2014visual], which robustly and reliably encode meaningful information. Such ensembles can also share units across semantically related inputs [yoshida2020natural].
We started our investigation by collecting and analysing representations from FF networks trained on Mnist, FashionMnist, Svhn and Cifar10, and defined, separately for each category, subsets of units (ensembles) that prominently and consistently activate in response to data in such category (subsection 4.2). These category-specific ensembles turn out to be composed of a few units, which is consistent with experimental findings on the sensory cortex [miller2014visual]. Furthermore, when image categories are characterised by a certain degree of visual similarity, the corresponding ensembles often share one or more units (Figure 3, Panel B). These results are consistent with phenomena observed in sensory cortices [yoshida2020natural] – as discussed in detail in subsection 2.2 – where ensembles are defined.
We then tested the ability of trained FF models to cope with new data, and we observed that the activations in response to an unseen input category formed in many cases a new ensemble, with similar characteristics of sparsity as the ones formed for other classes during training (Figure 4, Panels A and B). Furthermore, we noticed that the ensembles of unseen categories can share units with the ensembles of semantically-related categories used as training data (Figure 4, Panel C). These findings suggests that FF can perform well in zero-shot classification tasks, which is particularly relevant in view of the importance of zero/few-shot learning in human and animal cognitive performances [lake2015human].
While absent in the BP model, the existence of ensembles composed by a few units is not unique to FF. It is indeed observed also in BP/FF (subsection 4.2), where the ensembles turns out to be of comparable size. However, despite their similarity, the FF and BP/FF models are profoundly different, as demonstrated by the different interplay between inhibition and excitation in these models (see Appendix, G). We observe in this regard that the excitatory/inhibitory (E/I) balance plays a key role in the stability of cortical networks and in brain dynamics [gerstner2002spiking, deco2014local].
The sparsity of representations has computational benefits for sensory processing. [olshausen2004sparse] underline that sparsity may be the optimal encoding strategy for neuronal networks, for it is energy efficient. Sparsity also increases the memory-storage capacity, and eases readout at subsequent processing layers. [babadi2014sparseness] show that sparse and expansive coding (i.e., from a lower dimensional sensory input space to a higher dimensional neural representation) reduces the intra-stimulus variability, maximises the inter-stimulus variability, and allows optimal and efficient readout of downstream neurons. This is the reason why sparse and expansive transformations are widespread in biology, e.g., in rodents [mombaerts1996visualizing] or flies [turner2008olfactory].
Overall, our findings – focused on the emerging properties of representations – corroborate the idea that Forward-Forward might be a better model than Backprop for learning in the cortex [hinton2022forward]. As shown in section 4, besides being a non biological learning rule, Backprop elicits non-sparse and less biologically plausible representations.
This work is a starting point for further explorations in the field of biologically plausible representations. Promising avenues for future research in this field encompass exploring model compression through pruning [blalock2020state], with the design of new strategies based on the relevance of the ensembles, as well as delving into the dynamic evolution of ensemble size and organisation throughout the training process. Such investigations hold the potential to shed light on the formation, evolving interactions, and persistence or replacement of ensembles. Furthermore, these findings can be cross-referenced with neurophysiological data to assess the biological plausibility of ensembles in dynamic learning contexts.
6 Limitations
The limitations of the present work could be addressed by applying similar analyses to more complex datasets and a variety of tasks. To scale to a more challenging dataset may require the replacement of a fully connected network with more suitable architectures trainable with the Forward-Forward protocol, e.g. the CNNs recently introduced in [Papachristodoulou_2024]. This approach could provide insights into how data characteristics influence sparsity levels and the resulting ensemble-like structures. Moreover, as highlighted in [yang2023theory], the choice of hyperparameters potentially affects representation sparsity, as well as the goodness function of choice. Future work should explore the effects of these factors to further address the limitations identified.
[…]
Appendix A Computational resources
Training and subsequent experiments were conducted on an NVIDIA DGX A100 system. The system is equipped with 8 NVIDIA A100 GPUs, interconnected by NVLink technology, two AMD EPYC 7742 64-core CPUs, 1TB of RAM, and a 3TB NVME storage configured in RAID-0. Each GPU is equipped with $6912$ CUDA cores, $432$ Tensor cores and $40$ GB of high-bandwidth memory.
Appendix B Data
The Mnist dataset consists of pictures of handwritten Arabic numerals, from 00 to $9$, each represented as a grayscale image of size $28\times 28$. FashionMnist has been designed as a drop-in replacement to Mnist, offering a more challenging classification task. It consists of ten classes of clothing items, still represented as grayscale images with a resolution of $28\times 28$. Both datasets provide $60000$ training and $10000$ test images, balanced in terms of per-class numerosity.
Svhn contains colored images of digits from house numbers, captured by Google StreetView. The images are composed of $32\times 32$ RGB-encoded pixels. This dataset is slightly larger than the previous two, as it contains $73257$ data-points in the training set and $26032$ in the test set.
The Svhn images have been cropped in order to center the digit of interest within the frame. However, the presence of adjacent digits and other distracting elements, that have been kept within the images, introduces an additional layer of complexity when compared to Mnist and FashionMnist, where the subjects are prominently displayed against a uniform black background. The CIFAR10 consists of $60000$ colored natural images categorized in $10$ balanced classes. The dataset is split in $50000$ training images and $10000$ test images. Each image, like Svhn has a resolution of $32\times 32$ for each channel. Compared to previous datasets, this is the most challenging one for a fully connected network. The dataset split employed is provided by the TorchVision framework [torchvision2016].
Appendix C Training details
All our models (FF, BP/FF and BP), on all datasets (Mnist, FashionMnist, Svhn and Cifar10), have been optimised using Adam [kingma2017adam] with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ implemented in PyTorch [paszke2019pytorch]. A hyperparameter search has been performed to achieve sufficient accuracy for each model across all datasets. Every model was trained using batches of size $1024$.
Table 4: Hyperparameters selected to train our models.
Model Mnist FashionMnist Svhn CIFAR10
Epochs $1200$ $100$ $1000$ $1000$
FF Learning rate $0.01$ $0.01$ $0.0001$ $0.0001$
Epochs $300$ $300$ $200$ $200$
BP/FF Learning rate $0.0001$ $0.0001$ $0.0001$ $0.0001$
Epochs $80$ $80$ $80$ $80$
BP Learning rate $0.0001$ $0.0001$ $0.0001$ $0.0001$
Appendix D Activation patterns in deeper layers
In subsection 4.2 we claimed that in FF and BP/FF the images of a given category activate consistently a small set of units that we named ensembles, that share similarities to what is observed in sensory cortices. We reported in Figure 1 the activation map for Layer 1 (the first hidden layer) of FF trained on the Mnist dataset, and observed that very sparse ensembles emerge. In this section we show, in a similar fashion, the representations for Layers 2 and 3 (Figure 5 and Figure 6, respectively). We found that high sparsity also for deeper layers of this specific network; a qualitatively similar conclusion is reached for FF models trained on FashionMnist, Svhn and Cifar10. In BP/FF models a similar sparsity levels are observed, with the exception of the last layer that turns out to be not-sparse in each of the datasets we considered (see Table 3).
Figure 5: Activation patterns in a Multi-Layer Perceptron trained with the Forward-Forward algorithm, on the Mnist dataset. The image represents the activation map for neurons in Layer 2 for all images, grouped by class. A blue dot in position $(x,y)$ indicates that neuron $x$ is activated by input $y$; colorscale represents the intensity of such activation (incorrectly classified samples have been removed). Horizontal bands mark different categories; dark blue vertical lines mark active neurons. Each input category activates consistently a specific sets of neurons (ensemble). The sparsity measured according with the definition provided in subsection 3.5 is $0.84$.
Figure 6: Activation reported as in Figure 5, for Layer 3. Notice that there are only few units that activates significantly and does not play a role in discriminating categories. The role of this layer, in this experiment seems, not related to the classification task. Despite the low number of active units, the sparsity level of the representation is lower than that of Layer 2 ($S=0.67$), due to the noise of the inactive units.
Appendix E Activation patterns in different models
In Figure 1 Panel C we show the activation patterns in Layer 1 of FF trained on Mnist. For the purpose of a qualitative comparison, we show here analogous patterns for BP/FF and BP (see Figure 7 and Figure 8).
Figure 7: Activation pattern in Layer 1 of the BP/FF model trained on the Mnist dataset. The sparsity measure is $0.89$, comparable with the correspondent first layer of the FF model, reported in Figure 1, Panel C.
Figure 8: Activation pattern in Layer 1 of the BP model trained on the Mnist dataset. The sparsity measure is $0.32$ (non-sparse representation), about $\frac{1}{3}$ of the sparsity level measured in the analogous experiment with FF and BP/FF.
Appendix F Further results on representations of unseen categories and their ensembles
We showed in subsection 4.4 that a FF model trained on the FashionMnist dataset – deprived of one category – can respond at test time to this unseen category with a valid ensemble (Figure 4).
We report here the results of similar experiments, removing one category at a time. It turns out that, in each of the ten possible cases (we performed a single run for each category), the representations of the unseen category form a valid ensemble; we show three examples in Figure 9, different from the example shown in (Figure 4). It is with this situation in mind that we refer to “the ensembles related to unseen categories”.
Figure 9: Ensembles elicited by the FF model trained on FashionMnist deprived of one category (we show three examples: Pullover, Coat and Ankle boot). We report for the three categories, the activation value of each neuron in the first hidden layer (Layer 1), averaged on all images of the unseen category. Neuron index on the $x$ axis; average activation on the $y$ axis. Blue dots indicate units that are considered active according to the method described in subsection 3.5.
When an unseen category forms a valid ensemble, it generally exhibits a high level of integration with the ensembles associated with the categories encountered during training. This integration implies that it can share common units with ensembles belonging to related categories. We show in Figure 10 how the ensembles of missing categories (same examples as in Figure 9) integrate – by sharing units – with the other ensembles.
Figure 10: Shared units between the ensembles of unseen categories and the ensembles of categories seen during training (stripes delimited by the red lines). The results for Pullover, Coat and Ankle boot are shown.
Overall, these result relates to biological neural networks [yoshida2020natural, yuste2015neuron], where ensembles appear to be the functional building block of brain representations even in the absence of known stimuli.
Appendix G Distribution of excitatory and inhibitory connections
As we observed in subsection 4.2 FF and BP/FF have comparable sparsity levels and –when they are defined, the ensembles have comparable sizes. The presence of sparse ensembles suggests that a strong inhibition mechanism is at work, leaving only a few neurons active for each data sample. Inhibition in these architectures is the result of an interplay between the sign/magnitude of the weights and the sign/magnitude of the biases. Therefore it is natural to ask the following question: “are the FF and BP/FF similar also in this interplay between weights and biases?”. We find that the answer is no: the FF and BP/FF are indeed two profoundly different models that create sparse representations and ensembles with different mechanisms.
To show this, we consider for each neuron $i$ in a layer with width $n$ the fraction of its positive weights w.r.t. the total number of its input connections ($\varrho_{i}^{+}$ in the following), and its bias $\beta_{i}$. From these neuron-level quantities we construct their layer averages: $\varrho^{+}=\frac{1}{n}\sum_{i}\varrho_{i}^{+}$ and $\beta=\frac{1}{n}\sum_{i}\beta_{i}$. The neuron’s weights are strongly imbalanced towards inhibition when $\varrho_{i}^{+}\approx 0$ and, viceversa, strongly imbalanced towards excitation when $\varrho_{i}^{+}\approx 1$; when its $\varrho_{i}^{+}\approx 0.5$ we will say that the neuron’s weights are almost perfectly balanced; similar considerations holds for the biases, where a large and negative $\beta_{i}$ means strong inhibition for the $i$-th unit.
Focusing on the second hidden layer, we observe macroscopic differences of the empirical distribution of $\varrho_{i}^{+}$ among all the three models (see 11), with 1) a dominance of positive weights in the case of FF, 2) a bimodal distribution of $\varrho_{i}^{+}$ in BP/FF, with two populations of unbalanced units in opposite directions, and 3) a unimodal and approximately balanced distribution for all the neurons in the BP model.
[…]
Therefore, although the weights of the FF model appear imbalanced towards excitation, the average bias $\beta$ is very large and negative, and this might explain why, for this model, we observe such high sparsity values. On the other hand, the BP/FF model, have substantially zero average bias (within very small fluctuations), therefore the inhibitory mechanism at work here depends does not rely on a negative bias, but on the weights configuration.
From these results, we conclude that not only the training objective (i.e. goodness-based vs. categorical cross-entropy minimisation), but the specific training protocols (FF vs BP/FF) are determinant in creating different a interplay between excitation and inhibition, even when similar sparsity levels are observed.
Figure 11: Distribution of $\varrho_{i}^{+}$ in Layer 2 (Mnist dataset). In FF, the distribution is imbalanced, with most of the population of neurons having $\approx 65-75\%$ of excitatory weights. In BP/FF the distribution is bimodal with two populations of neurons: one inmbalanced towards excitation (right mode) and the other towards inhibition (left mode). The BP model is almost perfectly balanced between excitation and inhibition.
The data shown in Figure 11 refer to the second hidden layer of models trained on Mnist. We report the statistics $\varrho^{+},\beta$ for all the combinations of models, layers and datasets in Table 5 and Table 6.
Table 5: Average fraction of positive weights ($\varrho^{+}$) for each combination of models, datasets and layers. Results are expressed as mean $\pm$ std. dev. over one single run.
Dataset Layer FF BP/FF BP
1 $0.661\pm 0.031$ $0.534\pm 0.028$ $0.486\pm 0.022$
Mnist 2 $0.688\pm 0.014$ $0.445\pm 0.099$ $0.523\pm 0.019$
3 $0.882\pm 0.078$ $0.535\pm 0.071$ $0.52\pm 0.021$
1 $0.457\pm 0.099$ $0.509\pm 0.018$ $0.491\pm 0.021$
FashionMnist 2 $0.602\pm 0.074$ $0.423\pm 0.065$ $0.52\pm 0.02$
3 $0.416\pm 0.191$ $0.433\pm 0.019$ $0.52\pm 0.02$
1 $0.487\pm 0.062$ $0.499\pm 0.009$ $0.499\pm 0.006$
Svhn 2 $0.521\pm 0.033$ $0.427\pm 0.042$ $0.504\pm 0.011$
3 $0.583\pm 0.096$ $0.422\pm 0.031$ $0.522\pm 0.023$
1 $0.493\pm 0.027$ $0.502\pm 0.011$ $0.501\pm 0.009$
CIFAR10 2 $0.489\pm 0.044$ $0.43\pm 0.038$ $0.513\pm 0.012$
3 $0.612\pm 0.046$ $0.408\pm 0.034$ $0.523\pm 0.02$
Table 6: Average bias ($\beta$) for each combination of models, datasets and layers. Results expressed as mean $\pm$ std. dev. over one single run.
Dataset Layer FF BP/FF BP
1 $-1.57\pm 2.029$ $-0.007\pm 0.019$ $0.001\pm 0.021$
Mnist 2 $-1.66\pm 1.534$ $-0.017\pm 0.015$ $0.003\pm 0.018$
3 $-0.313\pm 0.151$ $-0.071\pm 0.02$ $0.002\pm 0.018$
1 $-0.694\pm 0.706$ $-0.007\pm 0.018$ $0.003\pm 0.021$
FashionMnist 2 $-1.858\pm 0.489$ $-0.011\pm 0.018$ $0.003\pm 0.019$
3 $-1.437\pm 1.213$ $-0.028\pm 0.009$ $0.003\pm 0.019$
1 $-0.662\pm 0.219$ $-0.033\pm 0.012$ $-0.004\pm 0.013$
Svhn 2 $-0.959\pm 0.089$ $-0.005\pm 0.011$ $0.003\pm 0.012$
3 $-0.962\pm 0.271$ $-0.001\pm 0.016$ $0.003\pm 0.011$
1 $-0.402\pm 0.1$ $-0.022\pm 0.008$ $-0.001\pm 0.016$
CIFAR10 2 $-0.703\pm 0.132$ $-0.006\pm 0.01$ $0.003\pm 0.016$
3 $-0.873\pm 0.076$ $-0.004\pm 0.013$ $0.002\pm 0.013$